# The One RING [Uncaptioned image]: a Robotic Indoor Navigation Generalist

### v1

@@ 3.3 Architecture
With this rich dataset of expert trajectories for random embodiments, a deep, high-capacity architecture is essential to learn a robust policy. In this section, we introduce our model architecture, shown in Fig. 3. At each timestep, Ring uses $N$ RGB images (one per camera) and a language instruction $l$ to predict an action distribution over a discrete action space. To account for different dimensions, we pad the RGB observations to make them square and resize them to $256\times 256$ before feeding them to the model. Ring’s architecture, inspired by PoliFormer [62], consists of a Visual Encoder, a Goal Encoder, a Transformer State Encoder, and a Causal Transformer Decoder with a linear actor-critic head. The Visual and Goal Encoders are frozen pre-trained models (a ViT and a language model, respectively) that encode the RGB observations and instruction $l$ into visual and goal token embeddings. Projections of these embeddings, along with a special STATE token vector, are stacked along the token axis and processed by the multi-layer Transformer State Encoder, which summarizes the observations at each timestep as the state embedding corresponding to the STATE token. Finally, the Causal Transformer Decoder performs explicit memory modeling over time, producing the current belief by causal attention on the state embeddings stacked along the temporal axis. The linear actor-critic head further predicts action logits over the action space as well as a value estimate. We provide more details about the architecture in the Appendix 9.
Figure 3: Our Ring model architecture. It accepts visual observations and a language instruction as inputs and predicts an action to execute. At RL finetuning, Ring also predicts a value estimate.
@@ 3.4 Training Paradigm
Recently, SPOC [17] showed that training policies with Behavior Cloning on large-scale expert trajectories in simulation leads to policies that effectively generalize to the real world. FLaRe [24] further introduced a robust and scalable method for finetuning such pretrained policies with On-policy Reinforcement Learning. RL finetuning introduces error recovery behaviors and mitigates the compounding errors typically encountered in imitation learning, leading to a substantial performance boost. We adopt the same recipe of first pretraining our policy on expert trajectories collected from randomized embodiments (Sec. 3.2), followed by finetuning with on-policy RL using the randomized embodiments in the AI2-THOR simulator [29].
Large-scale imitation learning with random embodiments. We train our policy using the architecture outlined in Sec. 3.3 on the collected trajectories. At each time step, the linear actor-critic head in Causal Transformer Decoder predicts the action logits. The cross-entropy loss is computed between the logits $\pi^{t}$ and the expert action. At training, we use a batch size of 240 trajectories and each trajectory has a temporal context window of 100 steps. We train our model on 8$\times$ H100 GPUs (80 GB memory per GPU) using the AdamW optimizer with a learning rate of $2{\cdot}10^{-4}$ for 80k iterations.
RL finetuning with random embodiments. Following the training recipe in FLaRe [24], we further perform a large-scale RL finetuning using AllenAct [49] on the randomized embodiments in simulation. Our training includes 1M random embodiments, 50k procedurally generated ProcTHOR houses [12] with $\sim$40k annotated 3D objects [18]. RL finetuning is specifically important for the policy to learn to navigate a diverse set of embodiments through trial-and-error. In particular, as the Ring policy lacks explicit information about its embodiment, it must implicitly infer it, which requires extensive exploration and trial-and-error. We use DD-PPO with $64$ parallel environments and $128$ rollout steps across 4 machines (each with 8$\times$ H100 GPUs) using the AdamW optimizer with a learning rate of $2{\cdot}10^{-5}$ for 40M training steps. Following FLaRe, we turn off the entropy term in the PPO loss to avoid catastrophic forgetting. For fair comparison, we use the reward function of [12],
$r_{t}=max(0,min\Delta_{0:t-1}-\Delta_{t})+s_{t}-\rho,$ (1)
where $min\Delta_{0:t-1}$ denotes the minimum L2 distance between the agent and any target object up to time $t-1$, $\Delta_{t}$ is the most recent L2 distance, $s_{t}$ is a success reward, and $\rho$ denotes the step penalty of $0.01$ to encourage the policy to finish the task efficiently. The agent must issue Done to indicate that it has found the target object to get the success reward $s_{t}=10$, otherwise $s_{t}=0$.
@@ 4 Experiments
Model Loss Train Embodiment Benchmark Embodiment
@@
Table 2: Zero-shot Results. Ring shows 0-shot generalization on 4 unseen embodiments. Stretch RE-1 and LoCoBot are evaluated on the Chores-$\mathbb{S}$ [17] ObjectNav benchmark (Unitree A1 is evaluated on a different benchmark which accounts for the agent’s lower height). “Stretch” without further qualification refers to the 2-camera variation on an RE-1 platform, as in [17]. All previous methods drastically fail to generalize to embodiments other than their training. Gray* numbers are evaluated on the training embodiment, otherwise evaluated zero-shot on an unseen embodiment.
Our experiments show that Ring operates effectively across a wide range of embodiments, including actual robots (Stretch RE-1, LoCoBot, and Unitree Go1) and human evaluation with Navigation Assistants, despite being trained exclusively in simulation without any direct exposure to real robot embodiments. Our key results are:
  1. 1.
    Ring generalizes zero-shot to 4 truly unseen embodiments, despite never being trained on them, and achieves state-of-the-art performance across multiple benchmarks (Sec. 4.1).
  2. 2.
    Our policy, trained solely in simulation on randomized embodiments, transfers directly to the real-world, on 3 real robots and navigation assistants (human evaluation). (Sec. 4.2).
  3. 3.
    Ring can be easily adapted to embodiment-specialized policies with minimal finetuning. It achieves better performance on each specific robot (Sec. 4.3).
  4. 4.
    Ring shows embodiment-adaptive behavior, adjusting its strategies based on the agent’s body (Sec. 4.4).
  5. 5.
    We present ablation studies and explore finetuning with collision penalties to enable the policy to take more conservative actions (Sec. 4.5).
@@ 4.1 Ring generalizes zero-shot to unseen embodiments
In this section, we perform zero-shot evaluations of all policies on four robot embodiments: Stretch RE-1 (with 1 or 2 cameras), LoCoBot, and Unitree A1 in simulation.
Baselines. For our baselines, we selected prior works in both imitation learning (IL) and reinforcement learning (RL). Each baseline is trained on a specific embodiment and evaluated in a zero-shot setting on four different embodiments. SPOC [16] is a supervised IL baseline trained on shortest-path expert trajectories in AI2-THOR. PoliFormer [62] is a state-of-the-art transformer-based policy in object goal navigation, trained from scratch using RL. FLaRe [24] is a approach for efficient policy finetuning that combines IL and RL. Specifically, SPOC [17] is trained with IL on Stretch RE-1 using $100$k expert trajectories; SPOC-$2.3$M is trained on more expert trajectories; Poliformer [62] is trained from scratch on each embodiment individually over $300$M RL steps; and FLaRe [24] finetunes SPOC on Stretch RE-1 with an additional $20$M RL steps.
Experimental Details. Ring is first trained with IL on $1$M expert trajectories collected from randomized embodiments in simulation, followed by finetuning with RL for an additional $40$M steps on the randomized embodiments (examples shown in Fig. 1-A and Fig. 2). Note that all four target embodiments were unseen during training, and information on embodiment is not provided during evaluation. We evaluate on the navigation benchmark in Chores-$\mathbb{S}$ [17], a simulation benchmark for household robot with 200 tasks across 200 scenes. For Unitree A1, we create a new, similar benchmark with 200 tasks adjusted for the robot’s lower height to ensure that all targets are feasible.
Results. Table 2 presents the zero-shot evaluation of all policies across four embodiments. We compare Success Rate and Success Weighted by Episode Length (SEL [15]), a metric measuring efficiency. The results indicate that all single-embodiment baselines struggle to generalize effectively to new embodiments, with performance declining as embodiment differences increase. For example, SPOC trained on the Stretch RE-1 with two cameras shows a gradual decrease in performance as the evaluation embodiment diverges from left to right on the top row in Table 2. It is worst when evaluated on the Unitree A1, which has a substantial height difference. In contrast, Ring exhibits strong generalization across all embodiments, despite not being trained on any of them, achieving an average absolute improvement of $16.7\%$ in Success Rate. In some cases, it outperforms the baseline trained on the target embodiment: PoliFormer trained on LoCoBot ($61.5\rightarrow 68.5$) and Unitree A1 ($55.3\rightarrow 72.0$). This shows that Ring benefits from training across random embodiments at scale, leading to a more effective navigation policy which even outperforms some embodiment-specialized policies (more detail on real robot experiments in Appendix 6.1).
@@ 4.2 Ring transfers to real-world embodiments despite being purely trained in simulation
Robot Evaluation We zero-shot evaluate our policy on 3 unseen robots in a real-world apartment. All evaluations are performed directly in a large-scale apartment (Fig. 4), without any further adaptation or real-world-specific finetuning. We used the same evaluation set of $15$ tasks for LoCoBot [12, 10, 62] ($3$ different starting poses with $5$ different targets), and $18$ tasks for Stretch RE-1 [17, 62, 24] ($3$ different starting poses with $6$ different goal specifications), respectively. We create a new evaluation set for Unitree Go1 with 3 starting poses and 4 objects (toilet, sofa, TV, trashcan) positioned to accommodate the robot’s lower height, ensuring that the objects can be visible from its lower viewpoint.
@@
Table 3: Real-world Results. Ring transfers zero-shot to the real-world without any finetuning. Gray numbers are evaluated on same embodiment as their training. Ring achieves $78.9\%$ success rate on average across $4$ real-world robots.
Human Evaluation To further demonstrate our policy’s generalization capability to unseen embodiments, we evaluate it as a navigation assistant with humans as new, unseen embodiments. We asked 5 participants to navigate in a real-world kitchen area, following the policy’s output actions on their phones. Each individual has unique characteristics, including step size, height, rotation angles, and camera-holding posture. Each person navigates to three different objects (Mug, Apple, Houseplant), resulting in a total of $15$ trajectories. We compare Ring with FLaRe [24], which was trained exclusively on the Stretch RE-1. Tab. 4 shows that Ring consistently outperforms FLaRe across target objects and different participants. Fig. 5 shows two qualitative results achieved by Ring (more details about human evaluation Appendix 6.2).
@@
FLaRE [24] Stretch RE-1 [Uncaptioned image] 40.0%
Ring-Zero-Shot Ring-Random [Uncaptioned image] 73.3%
Table 4: Human Evaluation. Five individuals navigate to 3 different objects ([Uncaptioned image] Apple, [Uncaptioned image] Houseplant, [Uncaptioned image] Mug) following the policy’s output actions on their phones in a kitchen area (example trajectories in Fig. 5). Ring-Zero-Shot shows much better generalization to human embodiment than the FLaRE baseline trained on Stretch RE-1.
Figure 6: Ring has embodiment-adaptive behavior, adjusting its navigation strategy based on the embodiment. The quadruped robot (B), due to its lower height, walks under the bed, while the taller Stretch-RE1 robot (A) navigates around it. In (C), an agent with the same height as Stretch-RE1 but a lower camera position initially attempts to move under the bed, assuming a shorter height. After colliding, it adapts its strategy and navigates around the bed, similar to Stretch-RE1.
@@ 4.3 Ring can efficiently adapt to an embodiment-specialized policy with minimal finetuning
Although Ring is a universal policy that works zero-shot across a broad range of embodiments, there are often cases where an embodiment-specialized policy is needed to achieve the best performance. In this section, we demonstrate that Ring can be easily adapted to a robot-specialized policy with minimal fine-tuning, resulting in better performance on the target embodiment.
Baselines. We use FLaRe [24] as a baseline. It has shown successful adaptation to new tasks and embodiments. This baseline is pretrained on Stretch RE-1 and finetuned on each of the three embodiments using up to $20$M RL steps.
Implementation Details. We finetune Ring, pretrained on randomized embodiments, on each individual robot for up to $20$M RL steps, while keeping all hyperparameters consistent with FLaRe to ensure a fair comparison. Following FLaRe, we repurpose two actions RotateBase(±$6^{\circ}$) to TiltCamera(±$30^{\circ}$) to allow camera movements for LoCoBot. Note that this movement is not allowed during the zero-shot evaluation.
Results. Fig. 7 shows that Ring adapts efficiently to specific embodiments with minimal finetuning, leading to embodiment-specialized policies with even better performance. For LoCoBot and Unitree-A1, FLaRe’s performance remains lower than on Stretch RE-1, indicating that pretraining on one embodiment and finetuning across embodiments cannot achieve the best results. This highlights the need for a policy that can consistently adapt to any embodiment with lightweight finetuning.
Figure 7: Embodiment-Specialized Adaptation. Ring, pretrained on randomized embodiments, shows efficient adaptation to robot-specialized policies with minimal fine-tuning. Baseline performance on LoCoBot and Unitree-A1 remains lower as they are fine-tuned on a different embodiment than the one used in pretraining. In contrast, Ring policy achieves consistent performance across all 3 embodiments, highlighting its capability for robot-specialized adaptation.
4.4 Ring has embodiment-adaptive behavior.
The behavior of optimal navigation policy, $\pi^{*}_{\theta}(a_{t}\mid o_{t}^{e},)$, should be strongly shaped by the agent’s body. For instance, an agent with a narrow FOV must explore more to effectively perceive its surroundings compared to an agent with a wider FOV. A smaller agent can navigate through narrow hallways or under furniture, and a larger agent may need to take more conservative paths. A question arises: does the policy show the same navigation behaviors across different embodiments, or does it adjust its strategy accordingly?
Our qualitative results confirm that the policy has an embodiment-adaptive behavior. In Fig. 6-A,B, both Stretch RE-1 and Unitree A1 start from the same pose behind the bed. The quadruped robot directly moves under the bed because of its lower height, while the Stretch RE-1 bypasses it. We observe that Ring implicitly infers embodiment parameters from visual observations and transition dynamics, dynamically adjusting its navigation strategy accordingly. It does not have access to any privileged information about its current body.
Visual observation reveals parameters, such as camera specs and, in some cases, agent height. However, visual information alone can be insufficient, leading the agent to rely on collisions to infer body dimensions. In Fig. 6-C, the agent matches the Stretch RE-1’s height but has a camera positioned as low as the quadruped’s. Initially, it assumes a lower height and attempts to go under the bed, but after colliding, it adjusts to maneuver around the bed, similar to the Stretch RE-1. This embodiment-adaptive navigation strategy adjustment is an interesting emergent behavior that would not have been possible without training across the exhaustive space of embodiments at scale.
4.5 Ablation Studies
A More Powerful Pretrained Visual Encoder. The default vision encoder used in our policies is the pretrained SIGLIP-ViT-B/16. In this section, we examine the impact of using a more powerful visual encoder on Ring’s performance. We train Ring-Large using OpenAI’s ViT-L/14 336px CLIP model [38]. Table 5 compares the results, showing that a stronger visual encoder significantly improves zero-shot performance across all four embodiments (approximately $9\%$ improvement on average). A larger visual encoder is particularly beneficial in our policy, as the visual observations are highly varied due to randomized camera parameters. To ensure fair comparison with the baselines and because ViT-L/14 is more computationally demanding, we chose to use the ViT-B/16 encoder for our main experiments. We will release the training code for the community for those interested in training with the larger visual encoder.
Include collision penalty to take safer routes. As our agents have randomized body dimensions and are not explicitly informed about their embodiment, they may occasionally collide before understanding their correct body size. In this section, we demonstrate that adding a small collision penalty of $0.1$ to the reward function can reduce collisions rate by $50\%$ (lowering the collision rate CR from $7\%$ to $4\%$). The resulting policy is more conservative, regardless of embodiment size.
To quantify these results, we created a custom benchmark similar to Chores-$\mathbb{S}$ [17], consisting of $2{,}000$ random embodiments across $2{,}000$ scenes. We evaluate $2$ different versions of our policy on this benchmark, comparing metrics such as Success Rate, Success Weighted by Collision (SC), Collision Rate (CR), and Safe Episode (percentage of episodes without any collisions). As shown in Table 6, adding the collision penalty reduces the collision rate (CR) ($7.77\%$rightarrow $4.03\%$) as well as increases the percentage of trajectories without collisions ($46.90\%$rightarrow $60.57\%$).
Model Visual Encoder Benchmark Embodiment
Stretch Stretch (Nav) LoCoBot Unitree A1
Ring SIGLIP-ViT-B/16 76.0 74.0 66.5 72.0
Ring-Large ViT-L/14 336px CLIP 83.8 77.7 75.3 79.9

### v2

@@ 3.3 Architecture
With this rich dataset of expert trajectories for random embodiments, a deep, high-capacity architecture is essential to learn a robust policy (Fig. 3). At each timestep, Ring uses $N$ RGB images (one per camera) and a language instruction $l$ to predict an action distribution over a discrete action space. To account for different dimensions, we pad the RGB observations to square and resize to $256\times 256$ before feeding them to the model. Ring’s architecture, inspired by PoliFormer [13], consists of a Visual Encoder, a Goal Encoder, a Transformer State Encoder, and a Causal Transformer Decoder with a linear actor-critic head. The Visual and Goal Encoders are frozen pre-trained models (a ViT and a language model, respectively) that encode the RGB observations and instruction $l$ into visual and goal token embeddings. Projections of these embeddings, along with a special STATE token vector, are stacked along the token axis and processed by the multi-layer Transformer State Encoder, which summarizes the observations at each timestep as the state embedding corresponding to the STATE token. Finally, the Causal Transformer Decoder performs explicit memory modeling over time, producing the current belief by causal attention on the state embeddings stacked along the temporal axis. The linear actor-critic head further predicts action logits over the action space as well as a value estimate. (More details in Appendix 10)
@@ 3.4 Training paradigm
We adopt the training recipe of pretraining our policy on expert trajectories collected from randomized embodiments (Sec. 3.2), followed by finetuning with on-policy RL using the randomized embodiments in the AI2-THOR simulator [59].
Large-scale imitation learning with random embodiments: We train our policy using expert trajectories collected from 1M randomized embodiments across 50k procedurally generated ProcTHOR houses [15], containing approximately 40k annotated 3D objects [60]. At each time step, the linear actor-critic head in the Causal Transformer Decoder predicts action logits, and a cross-entropy loss is computed between the predicted logits $\pi^{t}$ and the expert action. We use a batch size of 240 trajectories, each with a temporal context window of 100 steps. Training is conducted on 8$\times$ H100 GPUs (80 GB each) using the AdamW optimizer with a learning rate of $2\cdot 10^{-4}$ for 80k iterations.
Large-scale RL finetuning with random embodiments: Following the training recipe in FLaRe [14], we perform large-scale RL fine-tuning using AllenAct [61] on randomized embodiments in simulation. This fine-tuning phase is critical for enabling the policy to learn through trial and error how to navigate diverse embodiments. We use DD-PPO with 64 parallel environments and 128 rollout steps across 4 machines (each with 8$\times$ H100 GPUs), training for 40M steps using the AdamW optimizer with a learning rate of $2{\cdot}10^{-5}$. As in FLaRe [14], we disable the entropy term in the PPO loss to prevent catastrophic forgetting. For fair comparison, we adopt the reward function from [15]: $r_{t}=\max\left(0,min\Delta_{0:t-1}-\Delta_{t}\right)+s_{t}-\rho$, where $\min\Delta_{0:t-1}$ is the minimum L2 distance between the agent and the target object up to time $t{-}1$, $\Delta_{t}$ is the current L2 distance, $s_{t}$ is a success reward, and $\rho=0.01$ is a step penalty encouraging task efficiency. The agent must explicitly issue Done to receive the success reward ($s_{t}=10$); otherwise, $s_{t}=0$.
@@ 4 Experiments
Model Loss Train Embodiment Benchmark Embodiment: Success (SEL)
@@
Table 2: Zero-shot Results. Ring shows zero-shot generalization to four unseen embodiments. Unless otherwise specified, “Stretch” refers to the two-camera variant of the RE-1 platform, used in[12]. All prior methods fail to generalize effectively to embodiments beyond those seen during training. Gray* numbers indicate evaluation on the training embodiment; all others reflect zero-shot performance on unseen embodiments.
Our experiments show that Ring operates effectively across a wide range of embodiments, including Stretch RE-1, LoCoBot, Unitree Go1, and RB-Y1, despite being trained exclusively in simulation without any direct exposure to real robot embodiments. Our key results are: 1) Ring generalizes zero-shot to 5 truly unseen embodiments, despite never being trained on them, and achieves state-of-the-art performance across multiple benchmarks (Sec. 4.1). 2) Our policy, trained solely in simulation on randomized embodiments, transfers directly to the real-world, on 5 real robots and as navigation assistants (Sec. 4.2). (Videos in supplementary) 3) Ring can be easily adapted to embodiment-specialized policies with minimal finetuning. It achieves better performance on each specific robot (Sec. 4.3). 4) Ring shows embodiment-adaptive behavior, adjusting its strategies based on the agent’s body (Sec. 4.4). 5) We perform a collision analysis showing that Ring remains as safe—and in some cases even safer—than embodiment-specific policies(Sec. 4.5).
@@ 4.1 Ring generalizes zero-shot to unseen embodiments
We perform zero-shot evaluations of all policies on four robot embodiments: Stretch RE-1 (with 1 or 2 cameras), LoCoBot, and Unitree A1 in simulation.
Baselines. We select prior works from both imitation learning (IL) and reinforcement learning (RL) for comparison. Each baseline is trained on a specific embodiment and evaluated in a zero-shot setting across four different embodiments. SPOC [62] is a supervised IL baseline trained on shortest-path expert trajectories in AI2-THOR. PoliFormer [13] is a state-of-the-art transformer-based policy for object goal navigation, trained from scratch using RL. FLaRe [14] combines IL and RL for efficient policy fine-tuning. All baselines use similar architectures and comparable data, except for SPOC (reason to include SPOC 2.3M). Specifically, SPOC [12] (SPOC-$2.3$M) is trained with IL on Stretch RE-1 using $100$k ($2.3$M) expert trajectories; Poliformer [13] is trained from scratch on each embodiment individually over $300$M RL steps (more than other baselines in terms of training frames); and FLaRe [14] finetunes SPOC on Stretch RE-1 with an additional $20$M RL steps.
Experimental details. Ring is first trained with IL on $1$M expert trajectories collected from randomized embodiments in simulation, followed by finetuning with RL for an additional $40$M steps on the randomized embodiments. Note that all four target embodiments were unseen during training. We evaluate on the navigation benchmark in Chores-$\mathbb{S}$ [12], a simulation benchmark for household robot with 200 tasks across 200 scenes. For Unitree A1, we create a new, similar benchmark with 200 tasks adjusted for the robot’s lower height to ensure that all targets are feasible.
Results. Tab. 2 presents the zero-shot evaluation of all policies across four embodiments. We compare Success Rate and Success Weighted by Episode Length (SEL [45]), a metric measuring efficiency. The results indicate that all single-embodiment baselines struggle to generalize effectively to new embodiments, with performance declining as embodiment differences increase. In contrast, Ring exhibits strong generalization across all embodiments, despite not being trained on any of them, achieving an average absolute improvement of $16.7\%$ in Success Rate. In some cases, it outperforms the baseline trained on the target embodiment: PoliFormer trained on LoCoBot ($61.5\rightarrow 68.5$) and Unitree A1 ($55.3\rightarrow 72.0$). These 2 more challenging embodiments (lower FoV, low camera placement) make RL from scratch less effective. Ring benefits from more efficient learning by training across random embodiments at scale with more diverse navigation behaviors.
@@ 4.2 Ring transfers to real-world embodiments despite being purely trained in simulation
Robot evaluation. We zero-shot evaluate our policy on 4 unseen robots in a multi-room real-world apartment (Fig.6) without any real-world-specific finetuning (Tab. 4). We use the same evaluation set of 15 tasks for LoCoBot[15, 63, 13] (3 start poses × 5 targets) and 18 tasks for Stretch RE-1 [12, 13, 14] (3 poses × 6 goals). For Unitree Go1, we create a new set with 3 start poses and 4 objects (toilet, sofa, TV, trashcan) placed to match its lower viewpoint. Ring matches or outperforms specialized policies, likely due to cross-embodiment (XE) training enabling robust sim-to-real transfer in the presence of real-world noise. We also deploy Ring on the RB-Y1 wheeled humanoid in an unstructured kitchen, where it successfully navigates to targets (trashcan, apple, houseplant, mug) at two different heights (standing/seated), using an iPhone 16 Pro camera mounted on the robot to stream visual observations (Fig. 1-B). (More detail in Appendix 7.1 & videos in supplementary.)
@@
Table 3: Real-world Results. Ring transfers zero-shot to the real-world without any finetuning. Gray numbers are evaluated on same training embodiment.
@@
FLaRE [14] Stretch RE-1 Refer to caption 40.0%
Ring-Zero-Shot Ring-Random Refer to caption 73.3%
Table 4: Human Evaluation. Five individuals navigate to 3 different objects ( Apple, Houseplant,Mug) following the policy’s output actions on their phones in a kitchen area (see Fig. 10).
Human evaluation. To further show Ring’s generalization to unseen embodiments, we evaluate our policy as a navigation assistant with humans as novel embodiments. Five participants navigated a real-world kitchen by following policy outputs on their phones. Each had unique characteristics (e.g., step size, height, rotation, camera posture) and was tasked with reaching three objects (Mug, Apple, Houseplant), yielding 15 trajectories. We compare Ring to FLaRe [14], trained only on Stretch RE-1. As shown in Tab.4, Ring consistently outperforms FLaRe across objects and users. Fig.10 shows two qualitative examples (more in Appendix 7.2).
Figure 4: Ring exhibits embodiment-adaptive behavior, adjusting its navigation strategy based on the robot’s physical configuration. The shorter quadruped robot (B) walks under the bed, while the taller Stretch-RE1 (A) navigates around it. In (C), an agent with the same height as Stretch-RE1 but a lower camera position initially attempts to go under the bed, mistakenly assuming it can fit. After a collision, it adapts and reroutes around the bed, similar to Stretch-RE1.
@@ 4.3 Ring can efficiently adapt to an embodiment-specialized policy with minimal finetuning
Figure 5: Embodiment-Specialized Adaptation. Ring, pretrained on randomized embodiments, adapts efficiently to individual robots with minimal fine-tuning.
Although Ring generalizes zero-shot across diverse embodiments, some scenarios benefit from embodiment-specialized policies for optimal performance. Here, we show that Ring can be easily adapted to a robot-specialized policy through minimal fine-tuning. Baselines. We compare with FLaRe [14], which demonstrates effective adaptation to new tasks and embodiments. It is pretrained on Stretch RE-1 and finetuned on each of the three test embodiments using up to $20$M RL steps. Implementation. We finetune Ring, pretrained on randomized embodiments, on each robot for up to $20$M RL steps, using the same hyperparameters as FLaRe for fair comparison. Following FLaRe, we repurpose RotateBase(±$6^{\circ}$) to TiltCamera(±$30^{\circ}$) for LoCoBot, enabling camera control not available during zero-shot evaluation. Results. As shown in Fig. 5, Ring adapts efficiently, achieving superior performance with minimal fine-tuning. For LoCoBot and Unitree-A1, FLaRe underperforms compared to Stretch RE-1, suggesting that pretraining on a single embodiment limits generalizability. This underscores the value of policies like Ring that can adapt quickly and consistently to new embodiments.
4.4 Ring changes its behavior across different embodiments
Ideally, an optimal policy would modify its behavior depending on the embodiment. For instance, a thinner robot can navigate through narrow hallways or under furniture, and a wider agent may need to take more conservative paths. Our qualitative results show that Ring exhibits embodiment-adaptive behavior. In Fig. 4-A,B, both Stretch RE-1 and Unitree A1 begin behind a bed. The low-profile quadruped moves directly under it, while Stretch RE-1 navigates around—demonstrating that Ring implicitly infers embodiment characteristics from visual input, without access to privileged body information. Visual input reveals cues like camera specs and, in some cases, the agent’s height. However, vision alone can be ambiguous, prompting the agent to rely on physical interactions—such as collisions—to refine its understanding. (Collision feedback may come from actual impacts or from sensors that anticipate collisions before they occur.) In Fig. 4-C, an agent with a low-mounted camera but tall body misjudges its own height, initially attempts to go under the bed, collides, and then reroutes like Stretch RE-1. This behavior is not present in the expert data but emerges during reinforcement learning through training across diverse embodiments. (Videos in supplementary)
Model Train Embodiment Safe Episode $\uparrow$
Stretch LoCoBot Unitree A1
PoliFormer [13] Stretch 45.0 - -
LoCoBot - 42.0 -
Unitree A1 - - 49.25
FLaRe [14] Stretch 62.0 - -
Ring-Zero-Shot Ring-Random 67.0 64.5 48.5
Table 5: Ring has more Safe Episodes (with no collisions) compared to embodiment-specific baselines.